How much pure acid must be added to 50mL of 25% acid solution to produce a mixture that is 60% acid ?
Answer by mananth(16949)

percent ---------------- quantity
Acid 25 ----------------50 ml
Pure acid 100 ----------------x ml
Mixture 60 ---------------- 50+xml

25*50+100x=60(50+x)
1250+100x=3000+60x
100x-60 x=3000 -1250
40 x = 1750
/ 40
x= 43.75 ml of Pure acid